Zhao wants $7,000 saved in 2 years to make a down payment on a house. How much money should she invest now at 6.1% compounded annually in order to meet her goal?

Answer by math_tutor2020(3816)


Answer: $6218.24


Work Shown:
A = 7000 dollars is needed two years from now
P = unknown deposit amount
r = 0.061 = decimal form of the 6.1% interest rate
n = 1 = to represent compounding annually
t = 2 = number of years

Solve for P using the compound interest formula
A = P*(1+r/n)^(n*t)
7000 = P*(1+0.061/1)^(1*2)
7000 = P*(1+0.061)^(2)
7000 = P*(1.061)^(2)
7000 = P*1.125721
7000 = 1.125721P
1.125721P = 7000
P = 7000/1.125721
P = 6218.23702320557 approximately
P = 6218.24

Check:
A = P*(1+r/n)^(n*t)
A = 6218.24*(1+0.061/1)^(1*2)
A = 7000.00335104
A = 7000.00
The answer is confirmed.